OpenStudy (anonymous):

Find the x-intercepts (if any) for the graph of the quadratic function. 6x2 + 12x + 5 = 0 Give your answers in exact form. Show your work.

OpenStudy (astrophysics):

You can use the quadratic formula \[x = \frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] where a = 6, b = 12, and c = 5